1. The dewhitening filter CH6 had no output. I disconnected the cable and put it to the monitor out of the AI filter.
So the dewhitening is not in the loop.
2. I have made a thermal control filter
BANK1: pole 0Hz, zero 1mHz / LF boost stage
BANK2: pole 1mHz, zero 30mHz / LPF stage
BANK3: pole 1Hz, zero 0.1Hz / phase compensation stage
It seems working with the gain of 0.05. As the thermal is very strong, the output has less than 10.
This means the we are effectively only using ~4bit. We need external filter.
Note that output of 30000counts were about 3V at CH6.
3. Measured End PZT feedback with and without the thermal control. The UGF seems to be 0.2Hz.
The suppression at 10mHz is ~100. This is so far OK.
I implemented a slow servo for green laser thermal control on c1scx.mdl. Ch6,7 of ADC and ch6 of DAC are assigned for this servo as below;
Ch6 of ADC: PDH error signal
CH7 of ADC: PZT feedback signal
CH6 of DAC: feedback signal to thermal of green laser
Note that old EPICS themal control cable is not hooked anymore.
I made a simple MEDM screen(...medm/c1scx/master/C1SCX_BCX_SLOW.adl) linked from GREEN medm screen (C1GCV.adl) on sitemap.
During this work, I noticed that some of the epics switch is not recovered by autoburt. What I noticed is filter switch of SUSPOS, SUSPIT, SUSYAW, SDSEN, and all coil output for ETMX.
I had no idea to fix them, probably Joe knows. I guess other suspensitons has the same problems.